Hello again online tutors! Here's my question:
If x-5, 2x, and 4x+6 are the first three terms of an arithmetic sequence, find x and the value of the three terms.
Thank you to anyone who can help!
Amelia B

Found 2 solutions by ikleyn, phillywily:
Answer by ikleyn(52787)

.
Amelia,

under this link

https://www.algebra.com/algebra/homework/Sequences-and-series/Sequences-and-series.faq.question.1110915.html

https://www.algebra.com/algebra/homework/Sequences-and-series/Sequences-and-series.faq.question.1110915.html


please find very similar problem that I solved couple of days ago at this forum.

It is a TWIN to your problem.

Consider it as your sample and then solve your problem in the same manner.

If you still have questions, post them to the forum or directly to me via the "Thank you" note.


Answer by phillywily(18)

Alright. Given that this is an arithmetic sequence, so the difference between the second term and the first term is equal to the difference between the third term and the second term.
Henceforth,
T3 - T2 = T2 - T1
4x+6-2x = 2x-(x-5)
2x+6 = x+5
X=-1
Now we substitute x into the equations.
The first term is x-5, hence
-1-5=-6
T2
= 2x
= 2(-1)
= -2
T3
= 4x+6
= 4(-1)+6
= -4+6
= 2
Hence, the three terms are -6, -2 and 2 respectively.